Consider an ordinary shower where hot water at 60 oC is mixed with cold water at 10 oC. If it is desired that a steady stream of water at 45 oC be supplied, determine the ratio of the mass flow rates of the hot to cold water. Assume the heat losses from the mixing chamber to be negligible and the mixing to take place at a pressure of 150 kPa.


1
Expert's answer
2021-12-16T00:10:02-0500

"E_{in} = E_{out}"

"m_ch_c+m_hh_h=(m_c+m_h)h_{final}"

"4.18\times283+x\times4.18(333) =(1+x)318\times 4.18"

"x = \frac{m_h}{m_c} =2.33"